OpenStudy (anonymous):

How do you find the derivative of xe^(lnx^2)?

OpenStudy (anonymous):

\[xe^{\ln (2x^2)}\]

myininaya (myininaya):

use the product rule then use chain rule

myininaya (myininaya):

first write y=x(2x^2)=2x^3 so all we really need now is the power rule if y is written this way

myininaya (myininaya):

y'=8x^2

OpenStudy (amistre64):

e^ln(x^2) = x^2 x(x^2) = x^3 y' = 3x^2 ... if I read it right :)

OpenStudy (amistre64):

myin is right for the other option ;)